# CASE – Condition-Aware Sentence Embeddings for Conditional Semantic Textual Similarity Measurement

### v2

@@ Abstract
Deshpande et al. (2023) proposed the C-STS task for measuring the similarity between two sentences under a given condition. They created a human-annotated dataset containing 18,908 instances where the semantic similarity between two sentences $s_{1}$ and $s_{2}$ is rated under two conditions $c_{\rm high}$ and $c_{\rm low}$ resulting in respectively high vs. low similarity between the two sentences. Moreover, they proposed cross-, bi- and tri-encoder baselines. Given a triplet $(s_{1},s_{2},c)$ a cross-encoder considers all interactions among tokens in $s_{1}$, $s_{2}$ and $c$. Although the cross-encoder setting benefits from having access to both $s_{1}$ and $s_{2}$ at the same time, it is computationally costly due to the large number of token interactions required for longer sentences and conditions. Moreover, it does not pre-compute conditional embeddings for the individual sentences, which is problematic when computing all pairwise similarities between sentences in a large set.
@@
Lin et al. (2024) proposed a tri-encoder-based C-STS method where they used routers and heavy-light attention Ainslie et al. (2023) to select the relevant tokens to a given condition. Specifically, they used the query vector of the [CLS] token of the condition to compute attention scores for the tokens in the sentence, which are subsequently used to compute a sentence embedding. Their method outperforms Hyper-CL for the tri-encoder setting.
Li et al. (2024) proposed a cross-encoder approach which predicts C-STS scores, without creating conditional embeddings. Prior work on C-STS has shown cross-encoders to perform poorly despite having access to both sentences simultaneously. Li et al. (2024) showed that this is due to the irrelevant information in the two sentences to the condition and proposed a token re-weighting strategy, inspired by object detection in computer vision Shi et al. (2023); Jaegle et al. (2021). Concretely, they compute two cross-attention matrices between $(s_{1},s_{2})$ and $c$, which are subsequently used to compute the correlations for the sentence or condition tokens. Although their method improves the performance of cross-encoder-based C-STS measurement, it still underperforms bi-encoders. Consequently, we propose bi-encoders for C-STS, which are computationally efficient and accurate.
@@
The LLM-based embeddings computed in § 3.1 has two main drawbacks. First, relative to MLMs-based sentence embeddings, LLMs produce much higher dimensional embeddings, which can be problematic due to their memory requirements (especially when operating on a limited GPU memory) and the computational cost involved in inner-product computations. In tasks such as dense retrieval, we must compare millions of documents against a query to find the nearest neighbours under strict latency requirements, and low-dimensional embeddings are preferable. Second, although LLMs are typically trained on massive text collections and instruction-tuned for diverse tasks Muennighoff et al. (2022), their performance on C-STS tasks have been poor Lin et al. (2024). As seen from our condition-aware prompt template, an LLM must be able to separately handle a variable condition statement and a fixed instruction. This setup is different from most tasks on which LLMs are typically trained, where the instruction remains fixed across all inputs. Therefore, it is important to perform a supervised fine-tuning step to LLM embeddings before they are used for C-STS tasks.
@@
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we apply six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional NV), SFR-Embedding-Mistral (4096 dimensional SFR), gte-Qwen2-7B-instruct (3584 dimensional GTE) and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional E5), sup-simcse-roberta-large (1024 dimensional SimCSE_large), and sup-simcse-bert-base-uncased (768 dimensional SimCSE_base). Further details of these models are provided in Appendix A.
We evaluate model performance on different pooling methods, prompt settings, and sentence constructions. Then, we evaluate the dimensionality reduction methods under supervised and unsupervised settings to learn the projection for CASE (PCA, ICA, and linear and non-linear MLPs). Both the linear and non-linear MLPs are Siamese bi-encoders with weight-sharing. As explained in § 3.2, they take $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ as the input embeddings, and return the projected embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$ as the outputs, which can be compared using a similarity metric such as the cosine similarity.
Our linear MLP is a single-layer linear transformation:
@@
Hyperparameters are tuned on a held-out validation set. We select a learning rate of $10^{-3}$ and a batch size of 512. The dropout rate is set to $20\%$ for linear MLP and $15\%$ for non-linear MLP.
To address annotation errors in the original C-STS dataset such as ambiguous and invalid conditions, Tu et al. (2024) re-annotated the original validation set. To conduct a more accurate and reliable evaluation, we use the original C-STS train set and re-annotated validation set. A 70-30 split is used for the re-annotated validation set, with randomly selected 70% of the data allocated for validation and the remaining 30% for testing. We use a single p3.24xl EC2 instance (8x V100 GPUs) for learning sentence embeddings, and a separate NVIDIA RTX A6000 GPU for supervised projection learning. Scikit-learn 1.3.0 is used for PCA and ICA. Torch 2.0.1+cu117 is used for MLP proejction. These settings are fixed across all experiments. In the case of reducing 4096-dimensional LLM-based sentence embeddings to 512-dimensional, training a linear MLP for CASE takes approximately 1.5 minutes, whereas a non-linear MLP requires about 5 minutes (wall-clock time).
@@
To generate CASE, we apply different ways to construct the prompt for LLM-based embeddings and to concatenate the condition and sentence for MLM-based embeddings. For LLM-based models, we have two main settings: (a) $\operatorname{cond}=f(c;I(s))$, where we encode the condition given the sentence, and (b) $\operatorname{sent}=f(s;I(c))$, where we encode the sentence given the condition as explained in § 3.1. For MLM-based models, we evaluate the two settings: (a) $\operatorname{CONC}(c+s)$, where we concatenate sentence after the condition, and (b) $\operatorname{CONC}(s+c)$, where we concatenate condition after the sentence. For each setting, we further evaluate the effect of subtracting the embedding for the condition, $\bm{c}=f(c;I(\emptyset))$.
@@
Table 3: Cosine similarity scores ($\cos$) for two sentences under different conditions are computed using NV embeddings (shown on the left side of the arrows) vs. using MLP projected $512$ embeddings (i.e. CASE embeddings given by (1)) (shown on the right side of the arrows). Compared to the unconditional similarity between the two sentences, conditional similarity scores computed using NV embeddings align well with the human ratings. Moreover, the conditional similarities are further appropriately amplified by the supervised projection.
@@
We compare the different sentence encoders in Table 2 for a fixed (i.e. 512) dimensional projection from their original embeddings using Spearman correlation. NV obtains the highest Spearman correlation across all sentence encoders and dimensionality reduction methods. GTE and SFR also achieve high Spearman correlation coefficients under the non-linear MLPs. SimCSE_large and SimCSE_base have the lowest Spearman correlation coefficients, regardless of the dimensionality reduction method being used. Overall, we see that LLM-based embeddings perform better than the MLM-based embeddings on the C-STS task.
@@
We propose CASE in a novel setting, which encodes the condition under sentence by subtracting the condition as a post-processing step when measuring C-STS. Our findings show that LLM-based sentence embeddings consistently outperform MLM-based embeddings for C-STS. Moreover, we introduce an efficient supervised projection learning method to reduce the dimensionality of LLM-based embeddings, while improving the performance in C-STS. CASE correctly compares the tokens relevant to a given condition between the two sentences as shown in further qualitative results in Appendix D.
@@
There is a large number of sentence encodes (over 240 models as at March 2025 evaluated on Massive Text Embedding Benchmark (MTEBMuennighoff et al. (2023) leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard). However, due to computational limitations, we had to select a subset covering the best performing (top ranked on MTEB at the time of writing) models for our evaluations. We will release our source code and the evaluation framework such that the research community can evaluate our proposed method with any sentence encoder for C-STS.
@@ Appendix B Isotropy for Embeddings
Another method to measure the isotropy is Principle Components (IPC) Mu et al. (2018). To address the numerical instability issues when using PCA for isotropy measurement in high-dimensional embedding spaces, we use an approximation method for computing the IPC. Given a normalized embedding matrix $E\in\mathbb{R}^{n\times d}$ where each row represents a unit vector, instead of computing the eigenvectors of the covariance matrix, we randomly sample $k$ unit vectors ${u_{1},u_{2},...,u_{k}}$ from the unit hypersphere $\mathbb{S}^{d-1}$. For each sampled direction $u_{i}$, we compute the function $F(u_{i})=\sum_{j=1}^{n}\exp(E_{j}\cdot u_{i})$, where $E_{j}$ represents the $j$-th embedding vector. The approximated IPC is then defined as the ratio between the minimum and maximum values of $F$: $\text{IPC}{\text{approx}}=\frac{\min{i\in{1,...,k}}F(u_{i})}{\max_{i\in{1,...,k}}F(u_{i})}$. If the approximated IPC value is close to 1, it means the embedding space shows high isotropy where vectors are uniformly distributed across all directions in the high-dimensional space instead of clustering in some dominant directions. Conversely, when the IPC value approaches 0, it means significant anisotropy in the embedding space. We use $k=1000$ random directions to compute the approximation of IPC. Table 5 shows that the post-processing step of subtracting the condition c gives higher approximated IPC values, indicating the improvement of isotropy.
@@ Appendix C Full Results for all Models
Each LLM encoder has its own preferred pooling method, recommended by the authors of those models. From Table 1 we see that the performance varies significantly depending on the pooling method being used, while the latent attention pooling used in NV reporting the best results. Note that NV does not support last or average pooling, while latent pooling is not supported by the other models.
@@
Appendix D Additional Qualitative Analysis
@@
We briefly show an example of our pipeline in Table 7. We treat the conditions as questions for sentences and extract the relevant information as answers to compare whether CASE can focus on the condition-related information. The overall similarity trend is in consistent with the actual ratings. For all sentence and answer pairs, the similarity scores increase after supervised projection. This demonstrates that the embeddings effectively capture the conditional meaning shift with higher condition-dependent similarity.

### v3

@@ Abstract
Deshpande et al. (2023) proposed the C-STS task for measuring the similarity between two sentences under a given condition. They created a human-annotated dataset containing 18,908 instances where the semantic similarity between two sentences $s_{1}$ and $s_{2}$ is rated under two conditions $c_{\rm high}$ and $c_{\rm low}$ resulting in respectively high vs. low similarity between the two sentences. Moreover, they proposed cross-, bi- and tri-encoder baselines. Given a triplet $(s_{1},s_{2},c)$ a cross-encoder considers all interactions among tokens in $s_{1}$, $s_{2}$ and $c$. Although the cross-encoder setting benefits from having access to both $s_{1}$ and $s_{2}$ at the same time, it is computationally costly due to the large number of token interactions required for longer sentences and conditions. Moreover, it does not pre-compute conditional embeddings for the individual sentences, which is problematic when computing all pairwise similarities between the sentences in a large set.
@@
Lin et al. (2024) proposed a tri-encoder-based C-STS method where they used routers and heavy-light attention Ainslie et al. (2023) to select the relevant tokens to a given condition. Specifically, they used the query vector of the [CLS] token of the condition to compute attention scores for the tokens in the sentence, which are subsequently used to compute a sentence embedding. Their method outperforms Hyper-CL for the tri-encoder setting. Liu et al. (2025) proposed a conditional contrastive learning method for C-STS, introducing weighted contrastive losses with a sample augmentation strategy. Although it improved performance for both bi- and tri-encoders, the former outperforms the latter.
Li et al. (2024a) proposed a cross-encoder approach which predicts C-STS scores, without creating conditional embeddings. Prior work on C-STS has shown cross-encoders to perform poorly despite having access to both sentences simultaneously. Li et al. (2024a) showed that this is due to the irrelevant information in the two sentences to the condition and proposed a token re-weighting strategy, inspired by object detection in computer vision Shi et al. (2023); Jaegle et al. (2021). Concretely, they compute two cross-attention matrices between $(s_{1},s_{2})$ and $c$, which are subsequently used to compute the correlations for the sentence or condition tokens. Although their method improves the performance of cross-encoder-based C-STS measurement, it still underperforms bi-encoders.
@@
The LLM-based embeddings computed in § 3.1 has two main drawbacks. First, relative to MLMs-based sentence embeddings, LLMs produce much higher dimensional embeddings, which can be problematic due to their memory requirements (especially when operating on a limited GPU memory) and the computational cost involved in inner-product computations. In tasks such as dense retrieval, we must compare millions of documents against a query to find the nearest neighbours under strict latency requirements, and low-dimensional embeddings are preferable. Second, although LLMs are typically trained on massive text collections and instruction-tuned for diverse tasks Muennighoff et al. (2022), their performance on C-STS tasks have been poor Lin et al. (2024). As seen from our condition-aware prompt template, an LLM must be able to separately handle a variable condition statement and a fixed instruction. This setup is different from most tasks on which LLMs are typically trained on, where the instruction remains fixed across all inputs. Therefore, it is important to perform a supervised fine-tuning step to LLM embeddings before they are used for C-STS.
@@
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we apply six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional NV), SFR-Embedding-Mistral (4096 dimensional SFR), gte-Qwen2-7B-instruct (3584 dimensional GTE) and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional E5), sup-simcse-roberta-large (1024 dimensional SimCSE_large), and sup-simcse-bert-base-uncased (768 dimensional SimCSE_base). Further details provided in Appendix A.
We evaluate model performance on different pooling methods, prompt settings, and sentence constructions. Moreover, we evaluate the dimensionality reduction methods under supervised and unsupervised settings to learn the projection for CASE (PCA, ICA, and linear and non-linear MLPs). Both the linear and non-linear MLPs are Siamese bi-encoders with weight-sharing. As explained in § 3.2, they take $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ as the input embeddings, and return the projected embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$ as the outputs, which can be compared using a similarity metric such as cosine.
Linear MLP performs a linear transformation:
@@
Hyperparameters are tuned on a held-out validation set. We select a learning rate of $10^{-3}$ and a batch size of 512. The dropout rate is set to $20\%$ for the linear MLP and $15\%$ for non-linear MLP.
To address annotation errors in the original C-STS dataset such as ambiguous and invalid conditions, Tu et al. (2024) re-annotated the original validation set. To conduct a more accurate and reliable evaluation, we use the original C-STS train set and re-annotated validation set. A 70-30 split is used for the re-annotated validation set, with randomly selected 70% of the data allocated for validation and the remaining 30% for testing. We use a single p3.24xl EC2 instance (8x V100 GPUs) for learning sentence embeddings, and a separate NVIDIA RTX A6000 GPU for supervised projection learning. Scikit-learn 1.3.0 is used for PCA and ICA. Pytorch 2.0.1 with cuda 11.7 is used for MLP proejction. These settings are fixed across all experiments. For reducing 4096-dimensional LLM-based sentence embeddings to 512-dimensional, training a linear MLP for CASE takes approximately 1.5 minutes, whereas a non-linear MLP requires about 5 minutes (wall-clock time).
@@
To generate CASE, we apply different ways to construct the prompt for LLM-based embeddings and to concatenate the condition and sentence for MLM-based embeddings. For LLM-based models, we have two main settings: (a) $\operatorname{cond}=f(c;I(s))$, where we encode the condition given the sentence, and (b) $\operatorname{sent}=f(s;I(c))$, where we encode the sentence given the condition as explained in § 3.1. For MLM-based models, we evaluate the two settings: (a) $\operatorname{CONC}(c+s)$, where we concatenate sentence after the condition, and (b) $\operatorname{CONC}(s+c)$, where we concatenate condition after the sentence. For each setting, we evaluate the effect of subtracting the condition embedding, $\bm{c}=f(c;I(\emptyset))$.
@@
Table 3: Cosine similarity scores ($\cos$) for two sentences under different conditions are computed using NV embeddings (shown on the left side of the arrows) vs. using MLP projected $512$ embeddings (i.e. CASE embeddings given by (1)) (shown on the right side of the arrows). Compared to the unconditional similarity between the two sentences, conditional similarity scores computed using NV embeddings align well with the human ratings. Moreover, the conditional similarities are further appropriately amplified by the supervised projection. Further qualitative evaluations shown in Appendix E.
@@
We compare sentence encoders in Table 2 for a fixed (i.e. 512) dimensional projection from their original embeddings. NV obtains the highest Spearman coefficient across all sentence encoders and dimensionality reduction methods. GTE and SFR also achieve high Spearman coefficients under the non-linear MLPs. SimCSE_large and SimCSE_base have the lowest Spearman coefficients, regardless of the dimensionality reduction method being used. Overall, we see that LLM-based embeddings perform better than the MLM-based embeddings on the C-STS task. Note that direct comparison between our method and other C-STS approaches is challenging due to differences in the train/test sets used in prior work. Nevertheless, we present our results alongside previous methods in Appendix D for completeness, which shows its superior performance.
@@
We propose CASE, which encodes the condition under sentence by subtracting the condition in a post-processing step when measuring C-STS. Our findings show that LLM-based sentence embeddings consistently outperform MLM-based embeddings for C-STS. Moreover, we introduce an efficient supervised projection learning method to reduce the dimensionality of LLM-based embeddings, while improving the performance in C-STS.
@@
There is a large number of sentence encoders (over 240 models as at March 2025 evaluated on Massive Text Embedding Benchmark (MTEBMuennighoff et al. (2023) leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard). However, due to computational limitations, we had to select a subset covering the best performing (top ranked on MTEB at the time of writing) models for our evaluations. We will release our source code and the evaluation framework such that the research community can evaluate our proposed method with any sentence encoder for C-STS.
@@ Appendix B Isotropy for Embeddings
Another method to measure the isotropy is Principle Components (IPC) Mu et al. (2018). To address the numerical instability issues when using PCA for isotropy measurement in high-dimensional embedding spaces, we use an approximation method for computing the IPC.
Given a normalized embedding matrix $E\in\mathbb{R}^{n\times d}$ where each row represents a unit vector, instead of computing the eigenvectors of the covariance matrix, we randomly sample $k$ unit vectors $\bm{u}_{1},\bm{u}_{2},...,\bm{u}_{k}$ from the unit hypersphere $\mathbb{S}^{d-1}$. For each sampled direction $\bm{u}_{i}$, we compute the function $F(\bm{u}_{i})=\sum_{j=1}^{n}\exp(\bm{e}_{j}{}^{\top}\bm{u}_{i})$, where $\bm{e}_{j}$ represents the $j$-th embedding vector. The approximated IPC is then defined as the ratio between the minimum and maximum values of $F$, given by (10).
$\mathrm{IPC}{\mathrm{approx}}=\frac{\min_{i\in{1,...,k}}F(u_{i})}{\max_{i\in{1,...,k}}F(u_{i})}$ (10)
If the approximated IPC value is close to 1, it means that the embedding space shows high isotropy where vectors are uniformly distributed across all directions in the high-dimensional space instead of clustering in some dominant directions. Conversely, when the IPC value approaches 0, it means significant anisotropy in the embedding space. We use $k=1000$ random directions to compute the approximation of IPC. Table 5 shows that the post-processing step of subtracting the condition c gives higher approximated IPC values, indicating the improvement of isotropy.
@@ Appendix C Full Results for all Models
Each LLM encoder has its own preferred pooling method, recommended by the original authors of those models. From Table 6 we see that the performance varies significantly depending on the pooling method being used, while the latent attention pooling used in NV reporting the best results. Note that NV does not support last or average pooling, while latent pooling is not supported by the other models.
@@
Appendix D Comparisons against Prior work
In Table 7, we compare our proposed method (CASE) against the best performances reported by the previously published C-STS methods. We report the results directly from their original publications and do not reproduce those methods in our comparisons. Moreover, when there are multiple encoder settings (e.g. bi-encoder, tri-encoder or cross-encoder) considered in prior work, we report the result for the best setting. The methods shown in the first four rows in Table 7 (Deshpande et al. (2023), Liu et al. (2025), Li et al. (2024b) and Yoo et al. (2024)) use the same training, validation and test sets from Deshpande et al. (2023). The penultimate method (Tu et al. (2024)) re-annotates the validation dataset from Deshpande et al. (2023) and further split it into a train/test (70% vs. 30%). They then use this train split for training their method and evaluate the performance on their test split.
Our proposed CASE method uses the same re-annotated dataset by Tu et al. (2024), but with a different random split of 70-30 for train and test purposes because Tu et al. (2024) did not released their train/test random split. Due to the differences in encoder settings and train/test data splits used in prior work, direct comparisons of Spearman correlation coefficients is difficult. However, despite these discrepancies, we see that Tu et al. (2024) and CASE report significantly higher Spearman correlations compared to the rest of the methods.
Method Spearman
Deshpande et al. (2023) (bi-encoder) 47.5
Liu et al. (2025) (bi-encoder) 48.1
Li et al. (2024b) (cross-encoder) 43.8
Yoo et al. (2024) (tri-encoder) 39.6
Tu et al. (2024) (bi-encoder) 75.9
Our CASE (bi-encoder) 70.0
Table 7: Spearman correlation of previously proposed methods and ours (CASE). We report the best performing encoder setting (shown in brackets) from each published paper.
Appendix E Additional Qualitative Analysis
@@
We briefly show an example of our method in Table 8. We treat the conditions as questions for sentences and extract the relevant information as answers to compare whether CASE can focus on the condition-related information. The overall similarity trend is consistent with the actual ratings. For all sentence and answer pairs, the similarity scores increase after supervised projection. This demonstrates that the embeddings effectively capture the conditional meaning shift with higher condition-dependent similarity.